OpenStudy (anonymous):

What are the solutions of the quadratic equation below? -7x^2 - 23x + 10 = 0 Please help me.

OpenStudy (anonymous):

I know this, give me a second to type it in.

OpenStudy (aric200):

7x^2 - 23x +10=0 using the quadratic euation formula x=(-(-23) + sqrt((-23)*(-23) - (4*(-7)*10))/2*-7 and x=(-(-23) - sqrt((-23)*(-23) - (4*(-7)*10))/2*-7

OpenStudy (aric200):

x= 1/14(-23-sqrt(809)) or x= 1/14(sqrt(809)-23)

OpenStudy (anonymous):